# Mechanistic interpretability of large language models with applications to the financial services industry

### v1

@@ 4. Future Directions
We consider this paper to be the first step towards leveraging mechanistic interpretability techniques for understanding applications of LLMs in financial services. Some of the future research directions include: (i) Conducting experiments with more powerful open-source LLMs such as Mistral 7B and Llama-2 7B. This is where one may leverage fine-tuned open-source models from FinGPT project [12]. (ii) Working on more algorithmic examples of compliance financial tasks and preferably finding prompts that result in large logit differences. (iii) Perform path patching as a more sophisticated circuit discovery approach toward finding an end-to-end circuit for performing a compliance financial task.
@@ Acknowledgement
We would like to thank Dennis Lee (Chief IP Counsel & AGC Privacy & Security at DFS) and Sharon Greenbach (Sr. Counsel & Director, Regulatory Policy at DFS) for their helpful comments relevant to regulatory issues that arise in the financial industry. We also thank Callum McDougall, Shervin Minaee, and Christopher Olah for granting permission to use some of the figures and tables from their papers. Ashkan Golgoon benefited from stimulating discussions with Amirhossein Tajdini.

### v2

@@ 4. Future Directions
We consider this paper to be the first step towards leveraging mechanistic interpretability techniques for understanding applications of LLMs in financial services. Some of the future research directions include: (i) Conducting experiments with more powerful open-source LLMs such as Mistral 7B and Llama-2 7B. This is where one may leverage fine-tuned open-source models from FinGPT project [13]. (ii) Working on more algorithmic examples of compliance tasks in financial services and preferably finding prompts that result in large logit differences. (iii) Perform path patching as a more sophisticated circuit discovery approach toward finding an end-to-end circuit for performing a compliance financial task.
@@ Acknowledgement
We would like to thank Sharon O’Shea Greenbach (Sr. Counsel & Director, Regulatory Policy at DFS) and Dennis Lee (Chief IP Counsel & AGC Privacy & Security at DFS) for their helpful comments relevant to regulatory issues that arise in the financial industry and for carefully reviewing the manuscript. We also thank Callum McDougall, Shervin Minaee, and Christopher Olah for granting permission to use some of the figures and tables from their papers. Ashkan Golgoon benefited from stimulating discussions with Amirhossein Tajdini. The views and opinions expressed in this paper are solely our own and do not represent or reflect those of our employer.